OpenStudy (idealist10):

Find all curves y=y(x) such that the tangent to the curve at any point passes through a given point (x1, y1).

OpenStudy (idealist10):

@SithsAndGiggles

OpenStudy (anonymous):

For any point on the curve \((x_0,y_0)\), the slope of the tangent line to the curve is \(y'(x_0)\). The equation of the tangent line then is \[y-y_0=y'(x_0)(x-x_0)\] You know that the tangent line is to pass through a given point \((x_1,y_1)\), and so \[y_1-y_0=y'(x_0)(x_1-x_0)\] Since \(y=y(x)\) means \(y_0=y(x_0)\), you get an ODE that is linear in \(y(x_0)\). \[y_1-y(x_0)=y'(x_0)(x_1-x_0)\] or equivalently, \[y'(x_0)+\frac{1}{x_1-x_0}y(x_0)=\frac{y_1}{x_1-x_0}\]

OpenStudy (idealist10):

So I got dy/dx=(y1-y)/(x1-x), is this right? Do I solve for y from here?

OpenStudy (anonymous):

Yes if you replace \(y(x_0)\) with \(y\) you get the same ODE.

OpenStudy (idealist10):

So if dy/dx=(y1-y)/(x1-x), then how do you solve for y from here? \[\frac{ dy }{ y _{1}-y }=\frac{ dx }{ x ^{1}-x }\]?

OpenStudy (anonymous):

You can certainly treat this as a separable equation, so integrating both sides gives \[\int\frac{dy}{y_1-y}=-\ln|y_1-y|=\ln\frac{1}{|y_1-y|}\] and \[\int\frac{dx}{x_1-x}=-\ln|x_1-x|+C=\ln\frac{1}{|y_1-y|}+C\] Exponentiate both sides: \[\exp\left(\ln\frac{1}{y_1-y}\right)=\exp\left(\ln\frac{1}{|y_1-y|}+C\right)\] which gives \[\frac{1}{y_1-y}=\frac{C}{x_1-x}\]

OpenStudy (anonymous):

Alternatively, since the equation is linear in \(y(x_0)\), you can find the appropriate integrating factor. \[y'(x_0)+\frac{1}{x_1-x_0}y(x_0)=\frac{y_1}{x_1-x_0}\] suggests that the IF is \[\mu(x_0)=\exp\left(\int\frac{dx_0}{x_1-x_0}\right)=\frac{1}{x_1-x_0}\] This yields \[\begin{align*}\frac{1}{x_1-x_0}y'(x_0)+\frac{1}{(x_1-x_0)^2}y(x_0)&=\frac{y_1}{(x_1-x_0)^2}\\\\ \frac{d}{dx_0}\left[\frac{1}{x_1-x_0}y(x_0)\right]&=\frac{y_1}{(x_1-x_0)^2}\\\\ \frac{1}{x_1-x_0}y(x_0)&=\int\frac{y_1}{(x_1-x_0)^2}\,dx_0\\\\ \frac{1}{x_1-x_0}y(x_0)&=\frac{y_1}{x_1-x_0}+C \end{align*}\]

OpenStudy (idealist10):

So \[y _{1}-y=\frac{ x _{1}-x }{ c }\]

OpenStudy (idealist10):

Never mind. I got it! I got the right answer! Thank you!